An electron is projected vertically upward with a speed of 1.10×106m/s into a uniform magnetic field of 0.440 T that is directed horizontally away from the observer What is the radius of the motion?.
Answer #1

magnetic force acting on the electron,

F = q*v*B*sin(90)

m*a_rad = q*v*B

m*v^2/r = q*v*B

==> r = m*v/(B*q)

= 9.11*10^-31*1.1*10^6/(0.44*1.6*10^-19)

= 1.42*10^-5 m